OpenStudy (anonymous):

90 divided 14.4 don't know how to show my 5th grader

OpenStudy (tkhunny):

Have you considered moving the decimal? \(\dfrac{90}{14.4} = \dfrac{900}{144}\) Then, if you wanted, you could reduce the fraction. \(\dfrac{900}{144} = \dfrac{450}{72} = \dfrac{225}{36} = \dfrac{75}{12} = \dfrac{25}{4}\)

OpenStudy (anonymous):

No, we have to keep the decimal. I know the answer is 6.25 just don't know how to write out to figure out answer

OpenStudy (tkhunny):

The decimal doesn't go anywhere. You're just changing the numbers to make the calculation simpler. Every one of those fractions is equivalent to 6.25. Every one of them!

OpenStudy (anonymous):